# Dexterous Contact-Rich Manipulation via the Contact Trust Region

### v1

@@ X-E Hyperparameters for 2nd-order Dynamics Stabilization
TABLE IX: Hyperparameters for the MPC in Algorithm 3. No Heurisitcs and Closed-loop share the same parameters.
Figure 21: (a) Visualization of the reduced-order system for the Allegro in Section X-F1, which consists of four spheres. (b) Visualization of the IK result on the optimized solution, mapping the reduced-order system to the full system.
X-F Heuristics for Generating Allegro Contact Configurations
X-F1 A Reduced-Order Model
Inspired by [2, §5.4], instead of solving (40) directly on the full configuration space of ${q^{\mathrm{a}}}$, we first solve the problem on a reduced-order model, then map the solution to ${q^{\mathrm{a}}}$. For the Allegro hand, our reduced-order model consists of four spheres, each free to move in 3 dimensions with bounding-box joint limits (Figure 21a). Then, we solve for ${q^{\mathrm{a}}}$ by matching the fingertip positions to those of the spheres with Inverse Kinematics (IK), as shown in Figure 21b. Our IK procedure solves the following Quadratic Program (QP) iteratively,
@@ X-F1 A Reduced-Order Model
X-F2 Contact Sampling Distribution
@@
Acknowledgment
We thank Stephen Proulx and Velin Dimitrov for building the cube with active OptiTrack markers, and Giovanni Remigi for making the project website.

### v3

@@ A.6 Hyperparameters for 2nd-order Dynamics Stabilization
Table 12: Hyperparameters for the MPC in Algorithm 3. No Heuristics and Closed-loop share the same parameters.
Figure 22: Analysis of closed-loop tracking error as a function of the rollout horizon $H$. Performance is largely insensitive to this hyperparameter. For IiwaBimanual, there is a negligible correlation between $H$ and tracking error ($r=0.00082$). A minor negative correlation is observed for AllegroHand (r=0.059r=-0.059), which we attribute to violations of the quasi-dynamic assumption. Error bars represent $0.1\sigma$.
Figure 23: (a) Visualization of the reduced-order system for the Allegro in Section A.7.1, which consists of four spheres. (b) Visualization of the IK result on the optimized solution, mapping the reduced-order system to the full system.
A.7 Heuristics for Generating Allegro Contact Configurations
A.7.1 A Reduced-Order Model
Inspired by (Murray et al., 1994, §5.4), instead of solving (40) directly on the full configuration space of $q^{\mathrm{a}}$}, we first solve the problem on a reduced-order model, then map the solution to $q^{\mathrm{a}}$}. For the Allegro hand, our reduced-order model consists of four spheres, each free to move in 3 dimensions with bounding-box joint limits (Figure 23a). Then, we solve for $q^{\mathrm{a}}$} by matching the fingertip positions to those of the spheres with Inverse Kinematics (IK), as shown in Figure 23b. Our IK procedure solves the following Quadratic Program (QP) iteratively,
@@ A.7.1 A Reduced-Order Model
A.7.2 Contact Sampling Distribution
@@
(a) A goal which MPC with R-CTR successfully reached but the ETR baseline failed to reach. Note the initial contact points in frame (1) do not form a force-closure grasp. Without the dual feasibility constraint, ETR does not recognize the unilateral-ness of the initial contact configuration and fails catastrophically.
(b) A goal which both R-CTR and ETR-based MPC successfully reached. Note how the initial contact points form a nice anti-podal grasp on the cylinder.
Figure 24: Example trajectories for IiwaBimanual from running local MPC (Algorithm 2) to reach the goals generated in Section 5.1.1. While R-CTR-based MPC reaches both goals successfully, ETR-based MPC fails in (a) but succeeds in (b).
A.8 Local MPC Examples
For brevity and to maintain the narrative flow of the main text, we present detailed examples of local MPC (Algorithm 2) here, which include Figure 24 and Figure 25.
(a) A goal which MPC with R-CTR successfully reached but the ETR baseline failed to reach. The failure begins in frame (2), where a small gap appears between the thumb and the cube. Without the dual feasibility constraint, the ETR-based MPC incorrectly calculates that moving the thumb away from the object will reduce tracking error, leading to a catastrophic failure.
(b) A goal which both R-CTR and ETR-based MPC successfully reached. Note how the fingers form a nice closure grasp on the cube.
Figure 25: Example trajectories for AllegroHand from running local MPC (Algorithm 2) to reach the goals generated in Section 5.1.1. While R-CTR-based MPC reaches both goals successfully, ETR-based MPC fails in (a) but succeeds in (b).